2. Consider all samples of size 2 from this population:



2 5 6 8 10 12



a. Compute the mean () and variance () of the population.



b. List all samples of size 2 and compute the mean for each sample.



c. Construct the sampling distribution of the sample means.



d. Calculate the mean of the () of the sampling distribution of the sample means.



e. Calculate the variance of the () of the sampling distribution of the sample means.

1
Expert's answer
2022-05-26T23:52:10-0400

1.

 We have population values "2,5,6,8, 10, 12" population size "N=6"

"\mu=\dfrac{2+5+6+8+10+12}{6}=7.17"





"\sigma^2=\dfrac{1}{6}((2-7.17)^2+(5-7.17)^2+(6-7.17)^2""+(8-7.17)^2+(10-7.17)^2+(12-7.17)^2)=10.8"



2.

The number of possible samples which can be drawn without replacement is




"\dbinom{N}{n}=\dbinom{6}{2}=15"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,5 & 3.5 \\
 \hdashline
 2 & 2,6 & 4 \\
 \hdashline
 3 & 2,8 & 5\\
 \hdashline
 4 & 2,10 & 6 \\
 \hdashline
 5 & 2,12 & 7 \\
 \hdashline
 6 & 5,6 & 5.5 \\
 \hdashline
 7 & 5,8 & 6.5 \\
 \hdashline
 8 & 5,10 & 7.5 \\
 \hdashline
 9 & 5,12 & 8.5 \\
 \hdashline
 10 & 6,8 &7 \\
\hdashline
 11 & 6,10 & 8 \\
 \hdashline
 12 & 6,12 & 9 \\
 \hdashline
 13&8,10 & 9 \\
 \hdashline
 14 & 8,12 & 10 \\
 \hdashline
 15 & 10,12 &11 \\
 \hline
\end{array}"



3.

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 3.5 & 1 & 1/15 &0.23& 0.81 \\
 \hdashline
 & 4 & 1 & 1/15 & 0.26 & 1.06 \\
 \hdashline
 & 5 & 1 & 1/15 & 0.33& 1.66 \\
 \hdashline
 & 5.5 & 1 & 1/15 & 0.36& 2.02 \\
 \hdashline
 & 6 & 1 & 1/15 & 0.4 & 2.4 \\
 \hdashline
 & 6.5 & 1 & 1/15& 0.43 & 2.81 \\
 \hdashline
 & 7 & 2 & 2/15 & 0.93 & 1.86 \\
 \hdashline
 & 7.5 & 1 & 1/15 & 0.5 & 3.75 \\
 \hdashline
 & 8 & 1 & 1/15 & 0.53 & 4.26\\
 \hdashline
 & 8.5 & 1 & 1/15& 0.56& 4.82 \\
 \hdashline
 & 9& 2 & 2/15 & 1.2 & 10.8 \\
 \hdashline
 & 10 & 1 & 1/15 & 0.66 & 6.66 \\
 \hdashline
 & 11 & 1 & 1/15 & 0.73& 8.1 \\
 \hdashline
 Total & & 15 & 1 & 7.17 & 55.68 \\ \hline
\end{array}"




4.


"E(\bar{X})=\sum\bar{X}f(\bar{X})=7.17"


The mean of the sampling distribution of the sample means is equal to the mean of the population.



"E(\bar{X})=\mu_{\bar{X}}=7.17=\mu"



5.



"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"





"={55.68}-(7.17)^2=4.27"